# Improving Antibody Humanness Prediction using Patent Data

### v1

@@ 1 Introduction
Motivated by the importance of antibody humanness prediction, we present a Self-supervised learning framework designed to extract informative and contextually rich representations of amino acids and antibodies from Patented Antibody Database (SelfPAD). The Patented Antibody Database (PAD) is a rich and diverse source of information with functional characterization of antibodies, consisting of approximately $300,000$ unpaired sequences across seven different species (Krawczyk et al., 2021). Our approach to leverage the PAD for representation learning and immunogencity prediction departs from conventional approaches, pre-dominantly focused on leveraging information from Observed Antibody Space (OAS). The latter is the database with more than billion different antibody sequences sourced from more than $80$ different studies (Olsen et al., 2022a). What lacks in the OAS is functional characterization of antibodies which limits representation learning to self-supervised approaches focused on sequence completion tasks (Ruffolo et al., 2021; Olsen et al., 2022b; Shuai et al., 2023). In contrast, the PAD offers noisy association of sequences with function via more than $16,000$ different patent filings that can be translated into more than $1,000$ tasks, one for each set of potential targets (please see Section 4.1). This allows for multi-stage, multi-loss supervised learning, where a model is optimized with two different loss functions during the process. More specifically, we devote the initial stage to self-supervised representation learning to learn the representations of amino acids. We then switch the loss function to focus the final epochs entirely on a predictive task of interest, e.g. humanness prediction.
@@
We pose the initial stage focused on representation learning from patents as a weakly-supervised contrastive-learning problem, where each antibody sequence is associated with possibly multiple noisy identifiers of function and the objective is to learn a latent representation that groups them according to their patented objectives. In the first instance, this allows for learning of an antibody representation directly by the contrastive encoder that is informed pre-dominantly by the specific binding to different targets. One of our underlying assumptions is that the amino acid embeddings acquired through our proposed framework carry information about the functional and developability traits of antibodies as the patent filing is typically representative of the drug discovery process where the initial goal is to identify a potent binder, followed by subsequent improvements focused on refining its developability properties, encompassing factors such as immunogenicity, aggregation, viscosity etc. To demonstrate the utility of patent data for improving the immunogenicity prediction, we continue training the contrastive encoder to predict the humanness score of a given antibody sequence. We explore two different avenues for achieving transfer learning between the stages, the first via the pre-trained encoder itself and the second one via the learned embeddings of individual amino acids. During this stage of the training process, we also add a multi-layer perceptron (MLP) on top of the contrastive encoder and continue training on the patent data using the cross-entropy loss function. We evaluate the effectiveness of the learned immunogenicity model by performing inference on three different immunogenicity datasets published as part of prior work (see Sections  2 and  4.1). Our empirical analysis demonstrates that the final model learned from patent data achieves the new state-of-the-art performance on five out of six inference tasks, irrespective of the used metric. SelfPAD is available at: https://github.com/AstraZeneca/SelfPAD
@@ 2 Related Work
(b) Encoding sequence
Figure 1: SelfPAD Framework: a) The panel illustrates the training process that considers sequences associated with the same function identifier as positive pairs. b) The panel illustrates the architecture of the transformer-based sequence encoder.
@@
OASis is an approach quantifying the humanness of an antibody sequence based on peptide $k$-mers. It works by first segmenting a sequence into overlapping $9$-mer peptides and then querying these fragments against the Observed Antibody Space (OAS) (Olsen et al., 2022a) to estimate their prevalence within the human population. To comprehensively evaluate and compare humanness at the entire antibody level, it introduces the OASis identity score, which is calculated at the antibody level as the fraction of its $k$-mer peptides that surpass a user-defined prevalence threshold, where the threshold dictates the required fraction of the human population for a given peptide to be deemed human. OASis offers four predefined thresholds, capturing different stringency levels: i) loose ($\geq 1\%$ subjects), ii) relaxed ($\geq 10\%$ subjects), iii) medium ($\geq 50\%$ subjects), and iv) strict ($\geq 90\%$ subjects). For instance, at the relaxed threshold, a peptide is classified as human if found in at least $10\%$ of subjects. Accordingly, the OASis relaxed identity score for an antibody is computed as the fraction of its peptides found in at least $10\%$ of subjects. It is important to note that OASis may not perform effectively for antibodies with long CDR3 loops (Prihoda et al., 2022) since CDR3 loops are highly varied regions.
@@
When it comes to machine learning aspects of our work, it is related to setups involving self-supervised pre-training and task-specific fine-tuning. The paradigm of pre-training, whether supervised or unsupervised, and subsequent fine-tuning has proven to be a potent approach with promising results across various domains such as natural language processing (Mikolov et al., 2013; Pennington et al., 2014; Devlin et al., 2018; Liu et al., 2019; Conneau et al., 2019; Joulin et al., 2016; Collobert & Weston, 2008), computer vision and audio domains (Chen et al., 2020a; Grill et al., 2020; Oord et al., 2018; Caron et al., 2020; He et al., 2020; Falcon & Cho, 2020; Chen et al., 2020b). This progress was mainly enabled by taking advantage of spatial, semantic, or temporal structure in the data through data augmentation (Chen et al., 2020a), pretext task generation (Devlin et al., 2018) and using inductive biases through architectural choices (e.g., convolutional layers for images). In this particular instance, we leverage contrastive learning (Chen et al., 2020a) tailored for the multi-task setting at the initial epochs of the training process, and this stage can be considered a pre-training step. The second phase of the training process can be viewed as a form of fine-tuning. Unlike the conventional approach found in the literature, this fine-tuning occurs on the original dataset with the objective of learning to quantify the humanness of an antibody sequence. We should note that a well-curated patent dataset could facilitate the training of a multi-label, multi-task classifier. However, the identifiers of function currently available are noisy and may not be highly reliable. This is often due to a tendency toward risk aversion in safeguarding intellectual property, potentially resulting in patent filings containing sequences with varying functions. An alternative strategy for pre-training might involve disregarding the function and concentrating on self-supervised learning, which typically involves sequence completion tasks. However, we leave this as a potential avenue for future exploration. Finally, it’s important to emphasize that patent data serves as a unique information source for assessing the humanness of antibodies. Existing works, including those discussed here, predominantly rely on the OAS database. However, it is essential to note that the OAS database exhibits bias and lacks the necessary diversity to comprehensively characterize the entire population of human antibodies.
@@ 3 Method
Pre-training: The encoder consists of two components: i) a core transformer-based encoder that performs one-to-one mapping between amino acids from the antibody sequence and their context-dependent latent space embeddings and ii) a projection layer that maps the latent space representation of an antibody sequence to a fixed dimensional vector, independent of the sequence length. The transformer block has four layers and each one is implemented such that the layer normalization is applied prior to the self-attention and feedforward blocks (e.g., see Radford et al., 2019). We use a maximum sequence length of $180$ and the dimension of $32$ for the embedding and hidden layers. At the input, we first encode each amino acid in the sequence by using two trainable embedding layers, based on amino-acid type and position in the sequence, i.e., its positional encoding. We sum these two embedding vectors and concatenate the resulting embedding with additional $18$ features that are pre-computed based on the biophysical properties of amino acids such as their hydrophobicity, charge, volume etc. (please see Table 7 in the appendix for the full list of properties). We also inject a $\sigma$-scaled Gaussian noise (with $\sigma=0.01$) to the embeddings to learn a robust amino acid representation during training. We do not use any alignment for the sequences and do not explicitly condition the model with respect to the type of chain or the origin of species, i.e., we use each sequence as given. Since we use unigrams, the vocabulary size is twenty three to account for twenty amino acids as well as the tokens used for beginning and end of the sequence, and padding (#, \$, and * respectively). The projection layer includes a linear layer, followed by GELU activation and a final linear layer, where both linear layers have a dimension of $128$. We pre-trained the model with a batch size of $100$ for $1000$ epochs (see Figure 5 in the appendix for details).
@@
where $\psi(.)$ is the dot product similarity score, $\mathcal{L}_{c}$ is total contrastive loss, and $l(\bm{z}_{i1},\bm{z}_{i2})$ is the loss function for a positive pair $(\bm{z}_{i1},\bm{z}_{i2})$.
@@
Figure 2: After pre-trainnig: t-SNE visualisation of antibodies in the test set based on their a) Origin and b) Potential targets (showing only small subset of targets for better readibility). After fine-tuning the model to predict for humanness: c) t-SNE plot of heavy, kappa, lambda and vhh chains from different species. H-Heavy=Heavy chain from humans, O-Heavy: Heavy chain from other species.
Fine-tuning: We add a multi-layer perceptron (MLP) on top of the transformer-based encoder learned in the pre-training step. MLP consists of blocks with a residual connection, where each block has a linear layer with dimension $128\times 512$ followed by GELU activation, batch normalisation, dropout (with $p=0.3$) and another linear layer with dimension $512\times 128$. The final linear layer is paired with the softmax activation to give a binary classifer. To learn a robust representation of amino acids and antibodies, we also apply two types of noise during the training: i) shift noise: half of the time, we shift the starting position of a sequence by a randomly sampled number from the range $[0,10]$, and mask tokens are added to the beginning, shifting the sequence to the right. ii) swap noise: before inputting the resulting embedding into the MLP, we incorporate swap noise (Ucar et al., 2021), which involves randomly replacing selected embedding entries with values sampled from the same column.
We explored different strategies for the fine-tuning phase of the training process, where for example, one re-initializes the final four layers of encoder (effectively keeping only the learned amino acid embeddings and position encoding fixed) and fine-tunes by applying different learning rates to each layer of the contrastive encoder as well as the MLP blocks. More specifically, we increase the learning rate from low $0.0001$ to high $0.001$ as we move from lower to upper layers in the encoder. Alternatively, we also investigated the performance of the model that continues training of the contrastive encoder (i.e., no re-initialization of layers) on the humanness task during the fine-tuning epochs (including the extra MLP block). In the second phase of the training, we used cross-entropy loss with label smoothing (configured to $0.5$) and a batch size of $512$ for $25$ epochs.
@@ 4.1 Data Sources and Pre-processing
Patent data for antibodies can be a potentially valuable source of information for learning useful representations from their sequences. The primary purpose of patents is to provide legal protection rather than to serve a scientific knowledge, hence such database can be noisy. For example, although patented antibodies are tagged with potential targets, their relationship is not always clear. However, despite of its low signal-to-noise ratio, it can contain valuable information on relationships between family of antibody sequences and their potential targets. To illustrate its utility, we use Patented Antibody Database (PAD) (Krawczyk et al., 2021) that contains $16,526$ patent families from major jurisdictions such as US Patent and Trademark Office and World Intellectual Property Organization. These families hold $\sim 290k$ unique antibody chains (unpaired heavy and light chains as well as nanobodies) that are compiled in the PAD. Sequences in the database are manually tagged with potential target(s). If a sequence is associated with more than one target (e.g. ’il13’, ’il4’), we use the set of targets as the label (e.g. [’il13’, ’il4’]). We group sequences based on their association with potential targets. Sequences in each group constitute our positive samples for the corresponding target. We then randomly split the sequences into training ($\sim 260k$) and test sets ($\sim 29k$) by using stratified sampling to keep the proportion of antibodies from humans and non-humans same in both sets. Since the PAD also includes sequences from other publicly available datasets, we excluded $553$ therapeutics and $217$ immunogenicity data (Prihoda et al., 2022), along with $25$ humanization data (Marks et al., 2021), from the training set to avoid data leakage, considering that we utilize these datasets for evaluation purposes. Additionally, we excluded all sequences exhibiting at least $95\%$ sequence similarity with any sequence in the latter three datasets. The minimum Levenshtein distance of sequences in the three evaluation datasets from those in the training set of the PAD is presented in Table 1. Moreover, during pre-training, we drop any sequence that is not tagged with any target as well as any target that has less than 20 sequences in its group. After filtering, we end up with $1063$ potential targets and $\sim 136k$ sequences in the training set, which we then use to pre-train the encoder. In fine-tuning the pre-trained encoder for humanness prediction, we utilize all available training set sequences. The training set is then split into $90\%$ for training and $10\%$ for validation, and the model is fine-tuned with cross-entropy loss for $25$ epochs.
@@
Figure 3: Human versus Nonhuman sequences: PCA of embeddings from FT-SelfPAD for a) The test set of PAD, b) 553 therapeutics data from OASis, c) 217 immunogenicity data from OASis, d) 25 humanization data from Hu-mAb, where 25 therapeutic antibodies with immunogenicity problem (i.e. parental sequence) as well as their humanized versions, plotted together with rest of the sequences in the test set from the PAD (shown in blue).
553 Therapeutics is a dataset from Prihoda et al. (2022), also used in OASis, and originally sourced from IMGT mAb DB (Poiron et al., 2010). After removing four antibodies (letolizumab, lulizumab pegol, placulumab and glofitamab), we are left with $449$ paired sequences, consisting of $197$ human, $226$ humanized, $63$ chimeric, $41$ humanized/chimeric, $13$ murine, $6$ caninized, and $3$ felinized sequences. For humanness prediction evaluation, we classify the positive class as containing human and humanized sequences ($423$ sequences), while the rest are labeled negative.
217 immunogenicity refers to the dataset obtained from Prihoda et al. (2022) and also used in OASis. It contains reported ADA responses from $217$ paired sequences and is originally curated in Marks et al. (2021). We evaluate our model by scoring $217$ therapeutics for humanness. Sequences with immunogenicity $\geq 10$ are classified as nonhuman while the ones with $<10$ are as human.

### v3

@@ 1 Introduction
Motivated by the importance of antibody humanness prediction, we propose a weakly-supervised contrastive learning approach designed to extract informative and contextually rich representations of amino acids and antibodies from patented antibody database (PAD). As the contrastive learning is a well-known instance of self-supervised learning, we refer to our approach as SelfPAD. The patented antibody database is a rich and diverse source of information with functional characterization of antibodies, consisting of approximately $300,000$ unpaired sequences across seven different species (Krawczyk et al., 2021). Our approach to leverage the PAD for representation learning and immunogencity prediction departs from conventional approaches, pre-dominantly focused on leveraging information from Observed Antibody Space (OAS). The latter is the database with more than billion different antibody sequences sourced from more than $80$ different studies (Olsen et al., 2022a). What lacks in the OAS is functional characterization of antibodies which limits representation learning to self-supervised approaches focused on sequence completion tasks (Ruffolo et al., 2021; Olsen et al., 2022b; Shuai et al., 2023). In contrast, the PAD offers noisy association of sequences with function via more than $16,000$ different patent filings that can be translated into more than $1,000$ tasks, one for each set of potential targets (please see Section 4.1). This allows for multi-stage, multi-loss supervised learning, where a model is optimized with two different loss functions during the process. More specifically, we devote the initial stage to representation learning aimed at extracting an effective representations of amino acids. We then switch the loss function to focus the final epochs entirely on a predictive task of interest, e.g., humanness prediction.
@@
We pose the initial stage focused on representation learning from patents as a weakly-supervised contrastive-learning problem, where each antibody sequence is associated with possibly multiple noisy identifiers of function (i.e., potential targets) and the objective is to learn a latent representation that groups them according to their patented objectives. In the first instance, this allows for learning of an antibody representation directly by the contrastive encoder that is informed pre-dominantly by the specific binding to different targets. One of our underlying assumptions is that the amino acid embeddings acquired through our proposed framework carry information about the functional and developability traits of antibodies as the patent filing is typically representative of the drug discovery process where the initial goal is to identify a potent binder, followed by subsequent improvements focused on refining its developability properties, encompassing factors such as immunogenicity, aggregation, viscosity, etc. To demonstrate the utility of patent data for improving the immunogenicity prediction, we continue training the contrastive encoder to predict the humanness score of a given antibody sequence. We explore two different avenues for achieving transfer learning between the stages, the first via the pre-trained encoder itself and the second one via the learned embeddings of individual amino acids. During this stage of the training process, we also add a multi-layer perceptron (MLP) on top of the contrastive encoder and continue training on the patent data using the cross-entropy loss function. We evaluate the effectiveness of the learned immunogenicity model by performing inference on three different immunogenicity datasets published as part of prior work (see Sections  2 and  4.1). Our empirical analysis demonstrates that the final model learned from patent data achieves the new state-of-the-art performance on five out of six inference tasks, irrespective of the used metric.
@@ 2 Related Work
(b) Fine-tuning
Figure 1: SelfPAD Framework: a) Pre-training: The training process that considers sequences associated with the same potential targets as positive pairs. b) Fine-tuning: The architecture of the transformer-based sequence encoder used in pre-training as well as MLP, swap noise and shift noise used only during fine-tuning.
@@
OASis is an approach quantifying the humanness of an antibody sequence based on peptide $k$-mers. It works by first segmenting a sequence into overlapping $9$-mer peptides and then querying these fragments against the Observed Antibody Space (OAS) (Olsen et al., 2022a) to estimate their prevalence within the human population. To comprehensively evaluate and compare humanness at the entire antibody level, it introduces the OASis identity score, which is calculated at the antibody level as the fraction of its $k$-mer peptides that surpass a user-defined prevalence threshold, where the threshold dictates the required fraction of the human population for a given peptide to be deemed human. OASis offers four predefined thresholds, capturing different stringency levels: i) loose ($\geq 1\%$ subjects), ii) relaxed ($\geq 10\%$ subjects), iii) medium ($\geq 50\%$ subjects), and iv) strict ($\geq 90\%$ subjects). For instance, at the relaxed threshold, a peptide is classified as human if found in at least $10\%$ of subjects. Accordingly, the OASis relaxed identity score for an antibody is computed as the fraction of its peptides found in at least $10\%$ of subjects. It is important to note that OASis may not perform effectively for antibodies with long CDR3 loops (Prihoda et al., 2022), since these are highly varied regions.
@@
When it comes to machine learning aspects of our work, it is related to setups involving self-supervised pre-training and task-specific fine-tuning. The paradigm of pre-training, whether supervised or unsupervised, and subsequent fine-tuning has proven to be a potent approach with promising results across various domains such as natural language processing (Mikolov et al., 2013; Pennington et al., 2014; Devlin et al., 2018; Liu et al., 2019; Conneau et al., 2019; Joulin et al., 2016; Collobert & Weston, 2008), computer vision and audio domains (Chen et al., 2020a; Grill et al., 2020; Oord et al., 2018; Caron et al., 2020; He et al., 2020; Falcon & Cho, 2020; Chen et al., 2020b). This progress was mainly enabled by taking advantage of spatial, semantic, or temporal structure in the data through data augmentation (Chen et al., 2020a), pretext task generation (Devlin et al., 2018) and using inductive biases through architectural choices (e.g., convolutional layers for images). In this particular instance, we leverage contrastive learning (Chen et al., 2020a) tailored for the multi-task setting at the initial epochs of the training process, and this stage can be considered a pre-training step. The second phase of the training process can be viewed as a form of fine-tuning. Unlike the conventional approach found in the literature, this fine-tuning occurs on the original dataset with the objective of learning to quantify the humanness of an antibody sequence. We should note that a well-curated patent dataset could facilitate the training of a multi-label, multi-task classifier. However, the identifiers of function currently available are noisy and may not be highly reliable. This is often due to a tendency toward risk aversion in safeguarding intellectual property, potentially resulting in patent filings containing sequences with varying functions. An alternative strategy for pre-training might involve disregarding the function and concentrating on self-supervised learning, which typically involves sequence completion tasks. We do not pursue this, however, and leave it as a potential avenue for future exploration. Finally, it’s important to emphasize that patent data serves as a unique information source for assessing the humanness of antibodies. Existing works, including those discussed here, predominantly rely on the OAS database. However, it is essential to note that the OAS database exhibits bias and lacks the necessary diversity to comprehensively characterize the entire population of human antibodies.
@@ 3 Method
Pre-training. The encoder consists of two components: i) a core transformer-based encoder that performs one-to-one mapping between amino acids from the antibody sequence and their context-dependent latent space embeddings and ii) a projection layer that maps the latent space representation of an antibody sequence to a fixed dimensional vector, independent of the sequence length. The transformer block has four layers and each one is implemented such that the layer normalization is applied prior to the self-attention and feedforward blocks (e.g., see Radford et al., 2019). We use a maximum sequence length of $180$ and the dimension of $32$ for the embedding and hidden layers. At the input, we first encode each amino acid in the sequence by using two trainable embedding layers, based on amino-acid type and position in the sequence, i.e., its positional encoding. We sum these two embedding vectors and concatenate the resulting embedding with additional $18$ features that are pre-computed based on the biophysical properties of amino acids such as their hydrophobicity, charge, volume etc. (please see Table 8 in the appendix for the full list of properties). We also inject a $\sigma$-scaled Gaussian noise (with $\sigma=0.01$) to the embeddings to learn a robust amino acid representation during training. We do not use any alignment for the sequences and do not explicitly condition the model with respect to the type of chain or the origin of species, i.e., we use each sequence as given. Since we use unigrams, the vocabulary size is twenty three to account for twenty amino acids as well as the tokens used for beginning and end of the sequence, and padding (#, \$, and * respectively). The projection layer includes a linear layer, followed by GELU activation and another linear layer, where both of them have a dimension of $128$. We pre-trained the model with a batch size of $100$ for $1000$ epochs (see Figure 5 in the appendix).
@@
where $\psi(.)$ is the inner product similarity score, $\mathcal{L}_{c}$ is total contrastive loss, and $l(\bm{z}_{i1},\bm{z}_{i2})$ is the loss function for a positive pair $(\bm{z}_{i1},\bm{z}_{i2})$.
@@
Figure 2: After pre-training: t-SNE visualisation of antibodies in the test set based on their a) origin and b) potential targets (illustration for a small subset of targets). After fine-tuning the model to predict for humanness: c) t-SNE plot of heavy, kappa, lambda and VHH chains from different species (H denotes human and O heavy chains from other species).
Fine-tuning. For this stage, we add a multi-layer perceptron (MLP) on top of the transformer-based encoder learned in the pre-training step. MLP consists of blocks with a residual connection, where each one has a linear layer with dimension $128\times 512$ followed by GELU activation, batch normalisation, dropout (with $p=0.3$) and another linear layer with dimension $512\times 128$. The final linear layer is paired with the softmax activation to give a binary classifer. To learn a robust representation of amino acids and antibodies, we also apply two types of noise during training: i) shift noise: half of the time, we shift the starting position of a sequence by a randomly sampled number from range $[0,10]$, and add the mask tokens at the beginning. ii) swap noise: before inputting the resulting embedding into the MLP, we incorporate swap noise (Ucar et al., 2021), which involves randomly replacing selected embedding entries with values sampled from the same column.
We explored different strategies for the fine-tuning phase of the training process, where for example, one re-initializes the final four layers of encoder (effectively keeping only the learned amino acid embeddings and position encoding fixed) and fine-tunes by applying different learning rates to each layer of the contrastive encoder as well as the MLP blocks. More specifically, we increase the learning rate from low $0.0001$ to high $0.001$ as we move from lower to upper layers in the encoder. Alternatively, we also investigated the performance of the model that continues training of the contrastive encoder (i.e., no re-initialization of layers) on the humanness task during the fine-tuning epochs (including the extra MLP block). In the second phase of the training, we used cross-entropy loss with label smoothing (configured to $0.5$) and a batch size of $512$ for $25$ epochs. The code for SelfPAD is available at: https://github.com/AstraZeneca/SelfPAD
@@ 4.1 Data Sources and Pre-processing
Patent data for antibodies can be a potentially valuable source of information for learning useful representations from their sequences. The primary purpose of patents is to provide legal protection rather than to serve a scientific knowledge, hence such database can be noisy. For example, although patented antibodies are tagged with potential targets, their relationship is not always clear. However, despite of its low signal-to-noise ratio, it can contain valuable information on relationships between family of antibody sequences and their potential targets. To illustrate its utility, we use patented antibody database (PAD) (Krawczyk et al., 2021) that contains $16,526$ patent families from major jurisdictions such as US Patent and Trademark Office and World Intellectual Property Organization. These families hold $\sim 290k$ unique antibody chains (unpaired heavy and light chains as well as nanobodies) that are compiled in the PAD. Sequences in the database are manually tagged with potential target(s). If a sequence is associated with more than one target (e.g., ’il13’ and ’il4’), we use the set of targets as the label (e.g., {’il13’, ’il4’}). We group sequences based on their association with potential targets. Sequences in each group constitute our positive samples for the corresponding target. We then randomly split the sequences into training ($\sim 260k$) and test folds/sets ($\sim 29k$) by using stratified sampling to keep the proportion of antibodies from humans and non-humans the same in both sets. Since the PAD also includes sequences from other publicly available datasets, we excluded $553$ therapeutics and $217$ immunogenicity data (Prihoda et al., 2022), along with $25$ humanization data (Marks et al., 2021), from the training set to avoid data leakage, considering that we utilize these datasets for evaluation purposes. Additionally, we excluded all sequences exhibiting at least $95\%$ sequence similarity with any sequence in the latter three datasets. The minimum Levenshtein distance of sequences in the three evaluation datasets from those in the training set of the PAD is given in Table 1. During pre-training, we also drop any sequence that is not tagged with any target as well as any target that has less than 20 sequences in its group. After filtering, we end up with $1,063$ potential targets and $\sim 136k$ sequences in the training set, which we then use to pre-train the encoder. During fine-tuning of the pre-trained encoder for humanness prediction, we utilize all the available training set sequences. The training set is then split into two folds, $90\%$ for training and $10\%$ for validation, and the model is fine-tuned with cross-entropy loss for $25$ epochs.
@@
Figure 3: Human versus Nonhuman sequences: PCA of embeddings from FT-SelfPAD for a) the test set of PAD, b) 553 therapeutics data from OASis, c) 217 immunogenicity data from OASis, d) 25 humanization data from Hu-mAb, where 25 therapeutic antibodies with the immunogenicity problem (i.e., parental sequence) as well as their humanized versions were plotted together with other test fold sequences from the PAD (shown in blue).
553 Therapeutics is a dataset from Prihoda et al. (2022), also used in OASis, and originally sourced from IMGT mAb DB (Poiron et al., 2010). After removing four antibodies (letolizumab, lulizumab pegol, placulumab and glofitamab), we are left with $449$ paired sequences, consisting of $197$ human, $226$ humanized, $63$ chimeric, $41$ humanized/chimeric, $13$ murine, $6$ caninized, and $3$ felinized sequences. For humanness prediction evaluation, we classify the positive class as containing human and humanized sequences ($423$ sequences), while the others are labeled as negative.
217 immunogenicity refers to the dataset obtained from Prihoda et al. (2022) and also used in OASis. It contains reported ADA responses from $217$ paired sequences and it was originally curated in Marks et al. (2021). We evaluate our model by scoring $217$ therapeutics for humanness. Sequences with immunogenicity score $\geq 10$ are classified as nonhuman while the others are labelled as human.